Alan types a paper in 3hrs. Bill can type the same paper in 2hr. How long would it take them,working together,to type the paper?
Answer by nerdybill(7384)

Alan types a paper in 3hrs. Bill can type the same paper in 2hr. How long would it take them,working together,to type the paper?
.
Let x = time it takes for them working together
then
x(1/3 + 1/2) = 1
Multiplying both sides by 6 we get:
x(2 + 3) = 6
x(5) = 6
x = 6/5
or
x = 1 and 1/5 hours
or
x = 1 hour and 12 minutes